# The PLATO Mission

### v1

@@ 5.2 Support for the community
A second crucial activity within the ongoing PLATO-CS activity is to offer the community an electronic database containing a variability classification of the stars in the FoV prior to the launch of PLATO. An intense activity to train a meta-classifier by means of supervised and unsupervised classification is currently ongoing. This classification relies on machine-learning tools and is trained on Kepler light curves and class definition. Global and refined variability classifications will be done for the TESS all-sky survey [16]. With PLATO’s first FoV being recently fixed, these TESS variability classifications will be cross-matched with the sky position, and the content will be fine-tuned to be optimally suitable to help the community in its hunt for the best targets for all the science cases. The PLATO-CS variability catalogue will be made publicly available via a webform and will be continuously updated as newer results from the classifiers become available [17]. This variability catalogue will be a major source of information for GO applicants.
Other examples PLATO-CS can cover include, e.g., non-core exoplanet science as well as Solar System objects (see also examples in Rauer et al 254), such as microlensing events, stellar occultations by Solar System objects (e.g. asteroids), or studies of star-planet interaction mechanisms via phase curves detecting hot spots induced by orbiting planets [280].
We now provide three examples from stellar sciences which PLATO will be able to address.
5.3 Example 1: g-mode asteroseismology
@@
We highlight the major asset of PLATO-CS for the domain of gravity-mode (g-mode) asteroseismology, which got kick-started by the CoRoT mission with the discovery of period spacing patterns due to high-order g modes in the B-type dwarf HD 50230 [98]. While the 5-month CoRoT light curve was sufficient to discover the g-mode pattern of this slowly rotating pulsator, the g modes must be measured during at least two years to derive values for the internal rotation and mixing of rotating dwarfs. The real start of this science case thus came from the Kepler mission, whose 4-year light curves allowed for the development and application of g-mode asteroseismology along the main sequence. This meanwhile led to the internal rotation rates near the convective core in dwarfs covering the mass range from 1.3 to 9 M⊙ [see 5, for a summary]. PLATO-CS will be able to rely on light curves of 2-year duration, according to the current baseline observing scenario. This suffices to embark upon g-mode asteroseismology for a wide variety of masses, metallicities, and evolutionary stages, opening up asteroseismic probing of stars that will eventually explode as supernova as well as the most massive exoplanet host stars outside of PLATO’s core science. PLATO’s 2-year time base is critical to achieve this.
@@
The PLATO-CS potential for g-mode asteroseismology of exoplanet hosts is illustrated in blue in Fig. 16, which shows the Fourier transform in the form of a periodogram for the 684-d Kepler light curve of the Doradus pulsator KIC 8645874. This young intermediate-mass early-F type star has an effective temperature of 724090 K, , km s-1 and solar metallicity [295]. It was discovered to be a dipole prograde g-mode pulsator with an asteroseismic mass determination of M⊙. The star is found to rotate quasi-rigidly with a rotation period of d, which is about 5 times slower than the period of its dominant dipole g mode shown in Fig. 16 [304, 218].
@@
Figure 16 also shows the periodogram derived from this star’s 54-day TESS light curve in red. For comparison, we show in grey the periodogram resulting from a 54-day except of the Kepler light curve, to illustrate that the increased pixel size of TESS versus Kepler does not bring any contamination issues for this particular star, which has a visual magnitude of 9.92. The inset in Fig. 16 zooms in on two of the star’s g modes and reveals that the mode period precision downgrades tremendously by reducing the light curve from 684 d to 54 d (grey versus blue). It is also seen that the multiperiodic g-mode beating over the years and the limited frequency resolution of a 54-day light curve imply a different shape for the red and grey periodograms. Adding the TESS 54 d curve to the 684 d Kepler curve essentially gives the same result as the blue curve in Fig. 16 because the addition of the 2400 TESS data points to the 30620 Kepler data points, having a 2260 d gap in between, is not helpful to detect more modes in this particular example.
@@
The Kepler data of this star led to a derivation of the size and mass in its convective core due to internal mixing and allowed to deduce its evolutionary stage in terms of its central hydrogen mass fraction with a relative precision of [218]. PLATO’s 2-year light curves will deliver similar asteroseismic probing capacity to the blue curve in Fig. 16. This offers a good compromise between the 2-year duration of the light curve required to derive the internal rotation frequency along with the level of element mixing and the age, versus the number of suitable pulsators to perform g-mode asteroseismology by pointing at various positions in the sky.
5.4 Example 2: red giants and stellar ages
@@
Step-and-stare phases of duration shorter than two years are currently not foreseen in the nominal 4-year mission, but the instrument design makes it possible to take up such an observing strategy during the extended mission (see Section §7). Such step-and-stare phases offer limited capacity for the particular science case of g-mode asteroseismology of dwarfs, as the need to unravel the period spacing values with high precision requires the staring to cover at least one to two years [cf. Fig. 16 and 123].
@@
However, such a strategy would be optimal for other PLATO-CS topics, notably determining the age of red giant stars distributed across the Milky Way for galactic archaeology. This PLATO-CS science case is extensively discussed by Miglio et al [215], who showed that asteroseismic age-dating of red giants at level requires light curves with a duration of at least 150 d. However, the internal rotation of red giants can only be deduced from rotational splitting of their dipole mixed modes and requires a minimal duration of 2 years for the light curves, as also revealed from the detections of core rotation from Kepler data [32, 226].
The capacity to deduce the internal rotation and mixing properties of red giants and F-type dwarfs offers a critical tool to determine the age of PLATO’s core science exoplanet hosts at a level better than 10%. This is achieved by applying backtracking of the measured cumulative evolutionary effect via red giant models to the main sequence and by downscaling the models of Dor pulsators in mass [e.g., 219]. Such backtracking and downscaling of asteroseismically calibrated models will allow us to compute improved stellar models for the regime of the PLATO exoplanet hosts in the core science programme. This strategy will deliver models for the exoplanet hosts in the core science programme capable of age-dating at 10% level. The host stars themselves do not reveal oscillation modes capturing the internal rotation and mixing in the deep interior sufficiently well to achieve the required accuracy in age. For this reason, the PLATO core science calibration plan will rely on tens of thousands of red giants and thousands of Dor stars to be monitored during 2 years and covering proper ranges in mass, evolutionary stage, and metallicity. Of course, these red giants needed for the core science calibration then also serve for galactic archaeology as its requirements are less stringent, given that red giant age determination can already be done from d light curves.
5.5 Example 3: transient studies
@@
Outbursts in accreting compact binaries harbouring white dwarfs, neutron stars and black holes hold the potential to gain crucial information on disc instabilities and accretion processes. In particular low-mass X-ray binaries display outbursts precursors in optical wavebands prior to the X-ray detections and subsequent delays in optical to X-ray emission during the onset of outbursts [266, 129].
Outbursts from accreting white dwarf binaries of Cataclysmic Variable (CV) type have been studied for more than a century. These binaries have orbital periods ranging from 5 min to half a day and contain a white dwarf accreting material from a late type main sequence star or another white dwarf. Their outbursts arise through instabilities in the accretion disc and can have timescales from a few weeks to decades.
@@
Although there were only a dozen CVs in the Kepler field, the nature of the data allowed the accretion process to be studied in an unprecedented way. Dozens of normal and superoutbursts were observed from V344 Lyr which revealed positive and negative superhumps at different stages of the outburst cycle. These superhumps have periods slightly longer (positive) and shorter (negative) than the binary period and are due to a precessing and/or tilted accretion disc and whether the accretion stream impacts with the disc [288, 318, 240]. Kepler observations of V1504 Cyg and V344 Lyr also revealed that superoutbursts are preceded by a normal outburst giving insight to the physics of superoutbursts in general [62].
@@
Since then, many more CVs and accreting binaries have been studied using K2 and TESS. As one example of a CV observed using TESS in 2-min mode, we show in Fig. 17 two months of data of the CV VW Hyi. It shows a normal outburst preceding a superoutburst, with superhumps appearing at maximum brightness. The period of these positive superhumps starts to decrease as the outburst progresses. Observations such as these can be compared directly with models of superoutbursts [292, 175]. Other types of CVs have also been discovered to display unusual behaviour that went unnoticed using ground-based telescopes, such as the switch on and switch off of accretion in TW Pic [272] and fast optical enhancements in a number of magnetic CVs interpreted as micronova events [273].
@@
PLATO-CS will allow long term observations of relatively bright CVs, or systems which are usually faint but experience rare outbursts. More importantly, the potential exists to obtain high cadence, multi-colour photometry over the outburst cycle from PLATO´s fast cameras. This will allow for a search for variations in colour over orbital cycles and super-humps. Observations of VW Hyi made several decades ago [299] showed such colour variations can be present. PLATO-CS will also allow for the detection of longer period dwarf nova oscillations which can be used to probe the accretion process. We expect that outbursts from accreting binaries will be prime targets for pre-approved ToO observations.
@@ 6 PLATO Stellar Target Samples
6.1 Stellar Samples Definition
The PLATO mission has defined seven overall science objectives and four types of stellar samples (P1, P2, P4 and P5. P3 was dropped during mission development.). Here, we summarise the key parameters of the stellar samples (see Table 3, and Fig. 22 for a comparison to other missions).
@@ 6.1.1 The bright samples (P1 and P2)
These samples are at the core of the PLATO mission and consist of the brightest targets that PLATO will observe. Most of them will be part of the prime sample forming the main body of the final “PLATO Catalogue”. The sample P1 includes at least 15,000 dwarf and sub-giant stars (types F5 to K7), cumulative over the nominal mission, with mag and a noise level of 50 ppm in 1 h (see Table 3). We note that for the brighter stars in the sample ( mag) a noise level as low as 34 ppm in 1 hour can be reached (see Section 11). Sample P2 includes at least 1000 targets of the same type and noise performance with brightness . Targets in bright stellar samples observed by the normal cameras (N-CAM) are obtained with 25 sec sampling cadence.
A sub-sample of 300 stars will be observed with the two "fast" cameras (F-CAM) with 2.5 sec cadence, providing ’red’ and ’blue’ colour information. We note, however, that the F-CAMs cover only part of the central FoV of the normal cameras.
@@
| including: | 10 % | 50 | light curve | |||
| 5 % | 50 | centroids | ||||
| 9000 | 25 | imagettes |
@@ 6.1.3 The "statistical sample" (Sample P5)
6.1.4 The "prime" sample
A "prime sample" of up to 20000 targets will be selected out of the above stellar samples, supervised by the ESA PLATO Science Working Team (SWT), for observations with highest accuracy and will be treated with highest priority throughout the mission. Level-2 and Level-3 data products for the "prime sample" will form the core of the final "PLATO catalogue" consisting the final planet and stellar parameters derived by the mission.
The prime sample with targets in the first long pointing sky field will be defined nine months before launch and updated six months before every satellite sky field pointing. The PMC will organise ground-based follow-up observations to confirm the planetary candidates for this sample and measure the mass of the planets via its Ground-based Observing Programme (GOP).
The Science Management Plan grants the PMC proprietary rights over a small, pre-defined set of a maximum of 2000 targets in total over the 4 years of nominal mission duration. These targets will be selected using the first three months of PLATO observations of each field. From the Prime sample stars with brightness 11 for each sky field, the stars belonging to the lowest quartile (25%) of the noise distribution will be identified. The PMC proprietary targets will be 25% of these, with the condition that they will have a noise distribution similar to that of the original sample.
@@ 6.2 Light curves versus Imagettes
Let us have a look at the two examples to illustrate the capabilities of PLATO. The Table 5 assumes a homogeneous distribution of targets over the field of view and the maximum capabilities of the data processing system at camera level. In UC1 22,000 imagettes with 25 s sampling rate have been allocated per pointing. With two target fields, therefore, 44000 imagettes can be obtained. This number has to be compared to the requirements for the mission, which include at least 16000 imagettes for P1+P2, at least 9,000 imagettes in P5 and optionally additional imagettes for the P4 sample, adding to 30,000 required imagettes. The 44000 imagettes allocated here therefore include a substantial number of additional imagettes. This use case also foresees resources for 147,000 light curves per pointing with 600 s sampling. Assuming again 2 pointings over the course of the mission, this is much more than the required P5 sample size. In addition, resources have been devoted to light curves with higher sampling rates, centroids and housekeeping data, and calibration data. It is therefore an example where on-board light curve processing is maximised.
@@ 7 Target Fields and Observing Strategy
PLATO is a wide-field instrument whose field-of-view (FoV) covers 2132 deg2 on the sky (see Section 13). For comparison, the Kepler mission FoV was 115 deg2 [179], and the TESS satellite FoV covers 2300 deg2 [259], while CoRoT covered a square of 2.8x2.8 deg2 [18]. For illustration purposes, Fig. 18 shows approximate PLATO field positions on the sky in comparison to Kepler/K2 (K2 being the extended use of the Kepler satellite [152]) and CoRoT mission fields.
@@
PLATO will be launched into an orbit around the L2 Lagrangian point. The cruise and commissioning phases after launch can take up to 90 days. The duration of the nominal science operation phase of PLATO is then planned for 4 years. Extensions of the mission science operations are possible as the satellite has been designed with consumables for up to 8.5 years (see Section 12). For the design of the spacecraft, on-ground operations and mission performance studies, a set of baseline assumptions concerning the science observing strategy have to be made. The assumed baseline observing scenario therefore splits the 4 years science observation phase into 2 observing blocks of 2 years each, so-called long-pointings (LOPs). Alternative scenarios are possible, e.g. a longer field of 3 years duration followed by a phase of shorter step-and-stare pointings (SOPs). Another possible scenario would be, for example, to stare at one field only for the whole science operation phase. The spacecraft provides the technical flexibility to choose from such scenarios, and even adapt the strategy during the science operation phase if needed.
@@
The PLATO Input Catalogue (PIC) for first pointing of the satellite must be defined by ESA´s PLATO SWT at the latest 2 years before launch. An update is planned for 9 months before launch, including the definition of the so-called "prime sample" (see Section 6). The subsequent target fields and their "prime sample" members are defined 6 months before the start of each field. Once this mission is completed and in case an extended mission is granted, another evaluation for key regions which would deserve dedicated PLATO pointings can be made, unless extended observations of already covered fields are given priority.
@@
PLATO target sky regions are constrained by the mission science requirements defining the stellar samples to be observed (see Section 6) as well as technical constraints (e.g. Sun avoidance angles). In fact it turned out that due to the large field size and the requirement for long pointings, the possible sky target regions are constrained rather well. In a first step, stellar properties meeting the requirements of PLATO samples were investigated to produce an all-sky version of the PLATO Input Catalogue [PIC, 220]. In a next step, [229] presented the identification of possible sky regions for PLATO´s Long Pointings, taking into account technical boundary conditions, Gaia data and simulated signal-to-noise ratios for targets in these regions. First two provisional LOPs were presented, one for each hemisphere (Fig. 18).
@@
Meanwhile the final choice of two long pointing target fields has been made (Table 6), see Nascimbeni et al. in prep. for details on the selection and field properties. It turns out that both fields (north and south) meet the science requirements for the P1-P5 sample equally well. The choice of the first field was therefore made based on the availability of ground-based facilities for radial-velocity follow-up, in particular to provide masses for small planets. As a result of this assessment, the ESA PLATO SWT decided to start observations with the southern field (LOPS2).

### v2

@@ 5 The PLATO Complementary Science
First of all, PLATO-CS will offer the community an electronic database containing a variability classification and characterisation of all the stars brighter than about 13.5 mag in the TESS band and situated in PLATO’s fields. This database for the first long pointing (see Section 7) will be made available about a year prior to the launch. A meta-classifier is currently being trained for it, by means of supervised and unsupervised classification. It relies on machine-learning tools and is trained on Kepler and TESS light curves for the class definitions following Audenaert et al [17]. The PLATO-CS variability catalogue will be made publicly available and will be continuously updated as newer results from the classifiers become available [18]. This variability catalogue will be a major source of information for GO applicants (see Section 10).
Secondly, PLATO-CS actively extends the PLATO image and light curve simulator. This software tool takes into account all known instrumental noise sources to date to simulate realistic PLATO data [PlatoSim, 174]. PlatoSim is a versatile and publicly available suite of modules for the community to help assess PLATO’s capacity for a large variety of scientific goals. PLATO-CS also offers mock data simulated with PlatoSim. Along with Kepler and TESS light curves, the PLATO mock simulations serve as preparatory tools and testbeds to understand PLATO’s performance outside the core programme (Jannsen et al., in prep.).
Finally, PLATO-CS triggers the worldwide community to prepare competitive GO programs (see Section 10) on any topic not covered by the core science. PLATO-CS has predefined a series of such topics to ensure that these will in any case be the subject of GO applications [see 318, for an overview of these], including asteroseismology of massive and compact stars (Aerts et al., submitted), binarity and multiplicity, galactic structure of the Milky Way and Large Magellanic Cloud, transients and more broadly extragalactic science. However, many more topics are anticipated following ESA’s GO call. By means of illustrations, we now briefly discuss three examples defined by PLATO-CS as topics suitable for GO applications from the community.
5.1 PLATO-CS Example 1: gravity-mode asteroseismology
@@
Gravity-mode (g-mode) asteroseismology got kick-started by the CoRoT mission with the discovery of period spacing patterns due to high-order g modes in the B-type dwarf HD 50230 [103]. The 5-month CoRoT light curve of this star led to the first detection of a g-mode period spacing pattern of this slowly rotating pulsator. However, the real breakthrough in the derivation of the internal rotation and mixing inside rotating dwarfs came from the 4-years long Kepler light curves. These led to the near-core rotation rates of dwarfs covering the mass range from 1.3 to 9 M⊙ [see 5, for a summary]. With PLATO we have the opportunity to embark upon g-mode asteroseismology on a much larger scale, for a wide variety of masses, metallicities, and evolutionary stages, opening up asteroseismic probing of stars that will eventually explode as supernova as well as the most massive exoplanet host stars outside of PLATO’s core science.
@@
PLATO’s potential for g-mode asteroseismology is illustrated in in Fig. 16, which shows the Fourier transform in the form of a periodogram for the 684-d Kepler light curve of the Doradus pulsator KIC 8645874. This young early-F type star has an effective temperature of 724090 K, , km s-1 and solar metallicity [317]. It is a dipole prograde g-mode pulsator with an asteroseismic mass determination of M⊙. The star is found to rotate quasi-rigidly with a rotation period of d, which is about 5 times slower than the period of its dominant dipole g mode shown in Fig. 16 [328, 242].
@@
Figure 16 also shows the periodogram derived from this star’s 54-day TESS light curve in red and a 54-day string from the Kepler light curve in grey. This illustrates that the increased pixel size of TESS versus Kepler does not bring any contamination issues for this particular star, which has a visual magnitude of 9.92. The inset in Fig. 16 zooms in on two of the star’s g modes and reveals that the mode period precision downgrades tremendously by reducing the light curve from 684 d to 54 d (grey versus blue). It is also seen from the different shapes of the red and grey periodograms that the multiperiodic g-mode beating remains unresolved in a 54-day light curve.
@@
The Kepler data of this star led to the size and mass in its convective core and allowed to deduce its evolutionary stage in terms of its central hydrogen mass fraction with a relative precision of [242]. PLATO’s 2-year light curves can deliver similar powerful asteroseismic probing to the blue curve in Fig. 16, but now for tens of thousands of targets in the Milky Way.
5.2 PLATO-CS Example 2: galactic archaeology
@@
Step-and-stare phases of duration shorter than two years are currently not foreseen in the nominal 4-year mission, but the instrument design makes it possible to take up such an observing strategy during the extended mission (see Section 7). Such step-and-stare phases would be suitable for galactic archaeology. This science case is extensively discussed by Miglio et al [236], who showed that asteroseismic age-dating of red giants at level requires light curves with a duration of at least 150 d.
@@
The internal rotation of red giants, on the other hand, can only be deduced from rotational splitting of dipole mixed modes. This requires a minimal duration of 2 years for the light curves, as revealed from the detections of core rotation from Kepler data [32, 249]. PLATO offers this capacity from its onboard light curves, again for tens of thousands of red giants.
5.3 PLATO-CS Example 3: transient studies
@@
Outbursts in accreting compact binaries with white dwarfs, neutron stars or black holes hold the potential to gain crucial information on disc instabilities and accretion processes. In particular low-mass X-ray binaries display outburst precursors in optical wavebands prior to the X-ray detections and subsequent delays in optical to X-ray emission during the onset of outbursts [289, 140].
@@
Accreting white dwarf binaries of Cataclysmic Variable (CV) type have orbital periods ranging from 5 min to half a day and contain a white dwarf accreting material from a late type main sequence star or another white dwarf. Their outbursts arise through instabilities in the accretion disc and can have timescales from a few weeks to decades. The dozen CVs in the Kepler field allowed the accretion process to be studied in an unprecedented way. Both normal and superoutbursts were observed from V344 Lyr. These revealed positive and negative superhumps at different stages of the outburst cycle. These superhumps have periods slightly longer (positive) and shorter (negative) than the binary period and are due to a precessing and/or tilted accretion disc and accretion streams [310, 342, 260]. Kepler observations of V1504 Cyg and V344 Lyr also revealed that superoutbursts are preceded by a normal outburst giving insight to the physics of superoutbursts in general [68].
@@
Since then, many more CVs and accreting binaries have been studied using K2 and TESS - see the example of VW Hyi in Fig. 17. It shows a normal outburst preceding a superoutburst, with superhumps appearing at maximum brightness. The period of these positive superhumps decrease as the outburst progresses. Observations such as these can be compared directly with models of superoutbursts [314, 187]. Other types of CVs have also been discovered to display unusual behaviour that went unnoticed using ground-based telescopes, such as the switch on and switch off of accretion in TW Pic [295] and fast optical enhancements in a number of magnetic CVs interpreted as micronova events [296].
@@
PLATO will allow long term observations of relatively bright CVs, or systems which are usually faint but experience rare outbursts. More importantly, the potential exists to obtain high cadence, multi-colour photometry over the outburst cycle from PLATO´s fast cameras. This will allow for a search for variations in colour over orbital cycles and super-humps. Observations of VW Hyi made several decades ago [323] showed such colour variations can be present. PLATO-CS will also allow for the detection of longer period dwarf nova oscillations which can be used to probe the accretion process. We expect that outbursts from accreting binaries will be prime targets for ToO observations.
@@ 6 PLATO Stellar Target Samples
6.1 Stellar Samples Requirements
The PLATO mission has defined seven overall science objectives and four types of stellar samples (P1, P2, P4 and P5, P3 was dropped during mission development). Here, we summarise the key parameters of the stellar samples requirements set for the 4 year mission baseline (see Table 4, and Fig. 22 for a comparison to other missions). The actual target field is discussed in Section 7.
@@ 6.1.1 The bright samples (P1 and P2)
These samples are at the core of the PLATO mission and consist of the brightest targets that PLATO will observe. Most of them will be part of the prime sample forming the main body of the final “PLATO Catalogue”. The sample P1 includes at least 15,000 dwarf and sub-giant stars (types F5 to K7), cumulative over the nominal mission, with mag and a noise level of 50 ppm in 1 h (see Table 4). We note that for the brighter stars in the sample ( mag) a noise level as low as 34 ppm in 1 hour can be reached (see Section 11). Sample P2 includes at least 1000 targets of the same type and noise performance with brightness . The P2 targets are therefore a sub-set of the P1 sample, ensuring a certain number of very bright stars to be observed. Targets in these bright stellar samples observed by the normal cameras (N-CAM) are obtained with 25 sec sampling cadence.
A sample of 300 stars will be observed with the two "fast" cameras (F-CAM) with 2.5 sec cadence, providing ’red’ and ’blue’ colour information. We note, however, that the F-CAMs cover only part of the central FoV of the normal cameras.
@@
| including: | 10 % of the targets | 50 | light curve | |||
| 5 % of the targets | 50 | centroids | ||||
| 9000 targets | 25 | imagettes |
@@ 6.1.3 The "statistical sample" (Sample P5)
6.1.4 The "prime sample"
A "prime sample" of up to 20000 targets will be selected out of the above stellar samples (P1, P2 and P5), supervised by the ESA PLATO Science Working Team (SWT), for observations with highest accuracy and will be treated with highest priority throughout the mission. The PMC will organise ground-based follow-up observations to confirm the planetary candidates for this sample and measure the mass of the planets via its Ground-based Observing Programme (GOP). Level-2 and Level-3 data products for the "prime sample" will form the core of the final "PLATO catalogue" consisting the final planet and stellar parameters derived by the mission.
The prime sample with targets in the first long pointing sky field is to be defined at least nine months before launch and updated six months before every satellite sky field pointing. The SWT decided that 15000 prime sample targets should be selected in the first pointing field of PLATO (LOPS2, see Section 7), putting emphasize on the first field. The metrics to select the "prime sample" targets in the first field is currently under definition in the SWT. The criteria will be chosen such as to optimize the detectability of small planets with long orbital periods via transits in PLATO lightcurves as well as in the ground-based follow-up.
6.1.5 The propriety sample
The ESA Science Management Plan grants the PMC proprietary rights over a small, pre-defined set of a maximum of 2000 targets in total over the 4 years of nominal mission duration. These targets will be selected using the first three months of PLATO observations of each field. From the "prime sample" stars with brightness 11 for each sky field, the stars belonging to the lowest quartile (25%) of the noise distribution will be identified. The PMC proprietary targets will be 25% of these, with the condition that they will have a noise distribution similar to that of the original sample.
@@ 6.2 Light curves versus Imagettes
Let us have a look at the two examples to illustrate the capabilities of PLATO. In Table 6 we assume a homogeneous distribution of targets over the field of view and the maximum capabilities of the data processing system at camera level. In UC1 22,000 imagettes with 25 s sampling rate have been allocated per pointing. With two target fields, therefore, 44000 imagettes can be obtained. This number has to be compared to the requirements for the mission, which include at least 16000 imagettes for P1+P2, at least 9,000 imagettes in P5 and optionally additional imagettes for the P4 sample, adding to 30,000 required imagettes. The 44000 imagettes allocated here therefore include a substantial number of additional imagettes. This use case also foresees resources for 147,000 light curves per pointing with 600 s sampling. Assuming again 2 pointings over the course of the mission, this is much more than the required P5 sample size. In addition, resources have been devoted to light curves with higher sampling rates, centroids and housekeeping data, and calibration data. It is therefore an example where on-board light curve processing is maximised.
@@ 7 Target Fields and Observing Strategy
PLATO is a wide-field instrument whose field-of-view (FoV) covers 2132 deg2 on the sky (see Section 13). For comparison, the Kepler mission FoV was 115 deg2 [192], and the TESS satellite FoV covers 2300 deg2 [282], while CoRoT covered 8 deg2 [19]. For illustration purposes, Fig. 18 shows approximate PLATO field positions on the sky in comparison to Kepler/K2 (K2 being the extended use of the Kepler satellite [166]) and CoRoT mission fields.
@@
PLATO will be launched into an orbit around the L2 Lagrangian point. The cruise and commissioning phases after launch can take up to 90 days. The duration of the nominal science operation phase of PLATO is then planned for 4 years. Extensions of the mission science operations are possible as the satellite has been designed with consumables for up to 8.5 years (see Section 12). During long observing periods, the spacecraft has to be turned by 90 degrees every 3 months for sun protection reasons (see Section 12). For the design of the spacecraft, on-ground operations and mission performance studies, a set of baseline assumptions concerning the science observing strategy have to be made. The assumed baseline observing scenario therefore splits the 4 years science observation phase into 2 observing blocks of 2 years each, so-called long-duration Observing Phases (LOP). Alternative scenarios are possible, e.g. a LOP of 3 years duration followed by shorter Step-and-stare Observations Phases (SOP) with a minimum of 2 months each. Another possible scenario would be, for example, to stare at one field only for the whole science operation phase. The spacecraft provides the technical flexibility to choose from such scenarios, and even adapt the strategy during the science operation phase if needed.
@@
The PLATO Input Catalogue (PIC) for the first pointing of the satellite must be defined by ESA´s PLATO SWT at the latest 2 years before launch. An update is planned for 9 months before launch, including the definition of the so-called "prime sample" (see Section 6). The subsequent target fields and their "prime sample" members are defined 6 months before the start of each field. Once this mission is completed and in case an extended mission is granted, another evaluation for key regions which would deserve dedicated PLATO pointings can be made, unless extended observations of already covered fields are given priority.
@@
PLATO target sky regions are constrained by the mission science requirements defining the stellar samples to be observed (see Section 6) as well as technical constraints (e.g. Sun avoidance angles). In fact it turned out that due to the large field size and the requirement for long phase pointings, the possible sky target regions are constrained rather well. In a first step, stellar properties meeting the requirements of PLATO samples were investigated to produce an all-sky version of the PLATO Input Catalogue [asPIC, 243]. In a next step, [252] presented the identification of possible sky regions for PLATO´s LOP pointings, taking into account technical boundary conditions, Gaia data and simulated signal-to-noise ratios for targets in these regions. First two provisional LOPs (LOPN1 (north) and LOPS1 (south)) were presented, one for each hemisphere (Fig. 18).
@@
Meanwhile the final choice of two long pointing target fields has been made (Table 7), see Nascimbeni et al. 2024 (submitted) for details on the selection procedure and for the target field properties in terms of stellar populations and known transiting planets covered. A slight offset of 5 degrees of the southern field in comparison to the first choice was introduced to maximize pointing efficiency, leading to an updated southern field called LOPS2. It turns out that both fields (LOPN1 and LOPS2) meet the science requirements for the P1-P5 sample equally well. The choice of the first field was therefore made based on the availability of ground-based facilities for radial-velocity follow-up, in particular to provide masses for small planets. As a result of this assessment, the ESA PLATO SWT finally decided in June 2024 to start observations with the southern field (LOPS2).
The decision on whether to stay longer than the 2 years foreseen baseline on LOPS2 will be made by the ESA SWT after the first in flight science data have been analysed. Whether we move to the north after 2 years or later, the pointing coordinates of LOPN1 are unlikely to change significantly, because the large field of PLATO puts hard constraints. Potential SOP target fields, for the end of the baseline or for extended mission phases, have not been studied in detail yet. We note that significant work to prepare the respective input catalogues is needed before final decisions on such fields can be taken. Therefore, preparations for SOP fields have to start well in advance before observations can start. At the time of writing this manuscript, the SWT considers it unlikely that SOPs will be part of the nominal 4 year mission baseline. They are, however, kept as an option for mission extensions.